# Quantum-geometry-enabled Landau–Zener tunneling in singular flat bands

### v1

@@ Abstract
Here $\bm{\phi}=(\phi_{0},\phi_{1})^{\mathrm{T}}$ encodes the flat-band and dispersive-band components of the WS state, and $E$ is the WS eigenvalue. $E_{c}$ is given by Eq. (S3) and the flat band energy is zero. Eqs. (S23) and  (S24) resemble time-dependent Schrödinger equations, with $k_{y}$ playing the role of time.
@@
where $2k_{0}$ is the periodicity in $k$-space along the field direction [28]; for the present two-band model, $k_{0}=\pi$. Below we solve Eqs. (S23) and (S24) analytically in two limits: (i) an isolated-flat-band regime away from the BCP, and (ii) a strongly coupled regime near the BCP where LZT becomes essential.
@@
Isolated flat band limit.— For $k_{x}$ sufficiently far from the BCP and for weak fields, interband coupling is negligible and Eqs. (S23) and (S24) decouple [28]. Solving Eq. (S23) with the periodic boundary condition (S25) yields a WS ladder of the flat band (see Sec. III of the SM [53]):
@@
The WCC gives the average position of the Wannier function, displaced from the atomic orbitals by the Berry phase and thus determined by $d$. As shown in Fig. 1(c) for $d=0.8$, $\gamma(k_{x})$ exhibits an abrupt jump as $k_{x}$ crosses a BCP (gray dashed boxes). This reflects that each BCP carries a Berry phase $\xi\Delta\Phi_{\mathrm{B}}$, where $\Delta\Phi_{\mathrm{B}}=2\pi\sqrt{1-d^{2}}$ can take any value in $(0,2\pi)$ for $d\in(0,1)$ [50]. The closure of the WCC flow over the periodic BZ is ensured by paired BCPs of opposite chirality. Importantly, the abrupt jump at the BCP is merely an artifact of the isolated-band approximation: the interband Berry connections [Eq. (S8)] become non-negligible near the BCP. In the next section we show how the WS spectrum evolves through these regions once interband coupling is included.
@@
LZT near the singular point.— Close to the singular BCP ($|k_{x}|$ small), the isolated-flat-band approximation fails because the interband coupling is no longer negligible. We therefore compute the WS spectrum by diagonalizing the finite-size real-space 1D Hamiltonian (Sec. II A of the SM [53]), which is equivalent to solving Eqs. (S23) and (S24) with a truncated Fourier expansion in $k_{y}$. Because $k_{y}$ evolves periodically over the BZ, the WS spectrum can also be viewed as a Floquet quasienergy with periodicity $\Delta=eFa$ [56].
@@
To gain insights from quantum geometry, we analyze Eqs. (S23) and (S24) by treating intra- and interband Berry connections on an equal footing. For $k_{x}\to 0$, the Berry connections in Eqs. (S7) and (S8) are sharply peaked near $k_{y}=0$ with width $\sim|k_{x}|$. We therefore approximate the hybridization as occurring only at $k_{y}=0$ [28]. The net effect of Berry connections across $k_{y}=0$ can then be encoded in a unitary rotation relating $\bm{\phi}$ at $k_{y}=0^{-}$ and $0^{+}$:
@@
From Eqs. (S39) and (S32), $(\theta,\varphi)$ are functions of $F/t$ and $k_{x}$ and obey the scaling relations
@@
Once $(\theta,\varphi)$ are known, incorporating the dynamical phase (S32) and imposing the periodic boundary condition (S25) yields the WS quantization condition:
@@
Eq. (S44) reproduces Fig. 2(a–d), showing explicitly that $\theta$ controls the gap size and $\varphi$ controls the energy shift. If we neglect gap opening by setting $\theta=0$, Eqs. (11) and (12) apply with $\Phi_{\mathrm{B}}$ replaced by $\varphi$; the resulting dispersions are plotted as the orange dashed guides in Fig. 2(a–c).
@@
Using $(\theta,\varphi)$ for $d=1$ from Fig. 2(e) together with the scaling relation (S41), we obtain the WS spectra of the SFB from Eq. (S44) with $k_{0}=k_{0x}$ or $k_{0y}$ indicated in Fig. 3(a) (see Sec. VI of the SM [53]). For $F\parallel y$, the blue curves in Fig. 3(b) coincide with the two SFB ladders. For $F\parallel x$ [Fig. 3(c)], the slight discrepancy at small $k_{y}$ arises from additional LZT between the two dispersive bands near the Dirac cones at the $\mathrm{K}$ and $\mathrm{K}^{\prime}$ points, as indicated in Fig. 3(a).

### v2

@@ Abstract
Here $\bm{\phi}=(\phi_{0},\phi_{1})^{\mathrm{T}}$ encodes the flat-band and dispersive-band components of the WS state, and $E$ is the WS eigenvalue. $E_{c}$ is given by Eq. (5) and the flat band energy is zero. Eqs. (6) and  (7) resemble time-dependent Schrödinger equations, with $k_{y}$ playing the role of time.
@@
where $2k_{0}$ is the periodicity in $k$-space along the field direction [28]; for the present two-band model, $k_{0}=\pi$. Below we solve Eqs. (6) and (7) analytically in two limits: (i) an isolated-flat-band regime away from the BCP, and (ii) a strongly coupled regime near the BCP where LZT becomes essential.
@@
Isolated flat band limit.— For $k_{x}$ sufficiently far from the BCP and for weak fields, interband coupling is negligible and Eqs. (6) and (7) decouple [28]. Solving Eq. (6) with the periodic boundary condition (10) yields a WS ladder of the flat band (see Sec. III of the SM [53]):
@@
The WCC gives the average position of the Wannier function, displaced from the atomic orbitals by the Berry phase and thus determined by $d$. As shown in Fig. 1(c) for $d=0.8$, $\gamma(k_{x})$ exhibits an abrupt jump as $k_{x}$ crosses a BCP (gray dashed boxes). This reflects that each BCP carries a Berry phase $\xi\Delta\Phi_{\mathrm{B}}$, where $\Delta\Phi_{\mathrm{B}}=2\pi\sqrt{1-d^{2}}$ can take any value in $(0,2\pi)$ for $d\in(0,1)$ [50]. The closure of the WCC flow over the periodic BZ is ensured by paired BCPs of opposite chirality. Importantly, the abrupt jump at the BCP is merely an artifact of the isolated-band approximation: the interband Berry connections [Eq. (9)] become non-negligible near the BCP. In the next section we show how the WS spectrum evolves through these regions once interband coupling is included.
@@
LZT near the singular point.— Close to the singular BCP ($|k_{x}|$ small), the isolated-flat-band approximation fails because the interband coupling is no longer negligible. We therefore compute the WS spectrum by diagonalizing the finite-size real-space 1D Hamiltonian (Sec. II A of the SM [53]), which is equivalent to solving Eqs. (6) and (7) with a truncated Fourier expansion in $k_{y}$. Because $k_{y}$ evolves periodically over the BZ, the WS spectrum can also be viewed as a Floquet quasienergy with periodicity $\Delta=eFa$ [56].
@@
To gain insights from quantum geometry, we analyze Eqs. (6) and (7) by treating intra- and interband Berry connections on an equal footing. For $k_{x}\to 0$, the Berry connections in Eqs. (8) and (9) are sharply peaked near $k_{y}=0$ with width $\sim|k_{x}|$. We therefore approximate the hybridization as occurring only at $k_{y}=0$ [28]. The net effect of Berry connections across $k_{y}=0$ can then be encoded in a unitary rotation relating $\bm{\phi}$ at $k_{y}=0^{-}$ and $0^{+}$:
@@
From Eqs. (16) and (17), $(\theta,\varphi)$ are functions of $F/t$ and $k_{x}$ and obey the scaling relations
@@
Once $(\theta,\varphi)$ are known, incorporating the dynamical phase (17) and imposing the periodic boundary condition (10) yields the WS quantization condition:
@@
Eq. (19) reproduces Fig. 2(a–d), showing explicitly that $\theta$ controls the gap size and $\varphi$ controls the energy shift. If we neglect gap opening by setting $\theta=0$, Eqs. (11) and (12) apply with $\Phi_{\mathrm{B}}$ replaced by $\varphi$; the resulting dispersions are plotted as the orange dashed guides in Fig. 2(a–c).
@@
Using $(\theta,\varphi)$ for $d=1$ from Fig. 2(e) together with the scaling relation (18), we obtain the WS spectra of the SFB from Eq. (19) with $k_{0}=k_{0x}$ or $k_{0y}$ indicated in Fig. 3(a) (see Sec. VI of the SM [53]). For $F\parallel y$, the blue curves in Fig. 3(b) coincide with the two SFB ladders. For $F\parallel x$ [Fig. 3(c)], the slight discrepancy at small $k_{y}$ arises from additional LZT between the two dispersive bands near the Dirac cones at the $\mathrm{K}$ and $\mathrm{K}^{\prime}$ points, as indicated in Fig. 3(a).